# Adiabatic Electron Transfer in the Barrierless and Marcus-Inverted Regimes

### v3

@@ Abstract
The framework presented here highlights the importance of continuing to question the standard interpretation of the Marcus inverted region [23, 27]. As Marcus recalled in his Nobel lecture, experimental evidence for the inverted-region behavior was not obtained until "almost 25 years after it was predicted" but in that interim "experimentally in some reactions 100% formation of an electronically excited state of a reaction product has been observed" [28]. The possibility that the general difficulty observing the Marcus inverted region could be due to the adiabaticity of many reactions should be considered, whether results corresponding to the classically forbidden region or he region of extended barrierless transition ($\lambda<\Delta G^{\circ}<-\lambda_{\text{eff}}$) might explain a given experiment. Tunable Hamiltonian model systems, where both the driving force and electronic coupling can be systematically varied, may provide a route toward direct experimental testing of these predictions [29, 30, 31, 32, 30].

### v4

@@ Abstract
The framework presented here highlights the importance of continuing to question the standard interpretation of the Marcus inverted region [23, 27]. As Marcus recalled in his Nobel lecture, experimental evidence for the inverted-region behavior was not obtained until "almost 25 years after it was predicted" but in that interim "experimentally in some reactions 100% formation of an electronically excited state of a reaction product has been observed" [28]. The possibility that the general difficulty observing the Marcus inverted region could be due to the adiabaticity of many reactions should be considered, whether results corresponding to the classically forbidden region or the region of extended barrierless transition ($\lambda<\Delta G^{\circ}<-\lambda_{\text{eff}}$) might explain a given experiment. Tunable Hamiltonian model systems, where both the driving force and electronic coupling can be systematically varied, may provide a route toward direct experimental testing of these predictions [29, 30, 31, 32, 30].